OpenStudy (anonymous):

perform the indicated operation. x^2-2x-63/x^2-14x+49divided by x-9/x-7

OpenStudy (anonymous):

\[(x^2-2x-63)/(x^2-14x+49) divided by (x-9)/(x-7) = [(x-9)(x+7)/(x-7)(x-7)]*[(x-7)/(x-9)] = (x+7)/(x-7)\]

OpenStudy (anonymous):

thanks a lot!

OpenStudy (anonymous):

That didn't work out, anyway the answer is (x+7)/(x-7)

OpenStudy (anonymous):

thank you